OpenStudy (anonymous):

The explicit formula for the nth term of an arithmetic sequence is an = a1 + (n - 1) • d. What is the simple formula corresponding to the explicit formula if the first term of the sequence is -10 and the difference between terms in the sequence is 5? A. 5n - 5 B. 5n - 10 C. 5n + 15 D. 5n + 10 E. 5n - 15

OpenStudy (er.mohd.amir):

put a1=-10 and d=5

OpenStudy (anonymous):

5n-10

OpenStudy (er.mohd.amir):

-10+(n-1)5 =-10+5n-5 =5n-15

OpenStudy (anonymous):

ty